OpenStudy (laronjaxon):

Simplify x2-10x-24/x2-3x-108

OpenStudy (anonymous):

(x+2)/(x+9)

OpenStudy (laronjaxon):

Oh i didnt see this , can you and find the restrictions on the variable.

OpenStudy (anonymous):

x cant equal -9

OpenStudy (laronjaxon):

Thank you

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

\[x^{2} - 10x - 24 = x^{2} - 6x - 4x - 24\]\[x^{2} - 3x - 108 = x^{2} + 12x - 9x + 108\] \[\frac{ x^{2} - 6x - 4x - 24 }{ x^{2} + 12x -9x -108}\]\[=\frac{ (x-6)(x-4) }{ (x+12) (x-9) }\]